SEARCH FOR A VARIATION OF FUNDAMENTAL CONSTANTS

W. UBACHS, Department of Physics and Astronomy, VU University Amsterdam, De Boelelaan 1081, 1081 HV Amsterdam, Netherlands.

Since the days of Dirac scientists have speculated about the possibility that the laws of nature, and the fundamental constants appearing in those laws, are not rock-solid and eternal but may be subject to change in time or space. Such a scenario of evolving constants might provide an answer to the deepest puzzle of contemporary science, namely why the conditions in our local Universe allow for extreme complexity: the fine-tuning problem. In the past decade it has been established that spectral lines of atoms and molecules, which can currently be measured at ever-higher accuracies, form an ideal test ground for probing drifting constants. This has brought this subject from the realm of metaphysics to that of experimental science. In particular the spectra of molecules are sensitive for probing a variation of the proton-electron mass ratio μ, either on a cosmological time scale, or on a laboratory time scale. A comparison can be made between spectra of molecular hydrogen observed in the laboratory and at a high redshift (z=2-3), using the Very Large Telescope (Paranal, Chile) ^{a} and the Keck telescope (Hawaii) ^{b}. This puts a constraint on a varying mass ratio Δμ/μ at the 10^{−5} level. The optical work can also be extended to include CO molecules ^{c}. Further a novel direction will be discussed: it was discovered that molecules exhibiting hindered internal rotation have spectral lines in the radio-spectrum that are extremely sensitive to a varying proton-electron mass ratio. Such lines in the spectrum of methanol were recently observed with the radio-telescope in Effelsberg (Germany) ^{d}.

aF. van Weerdenburg, M.T. Murphy, A.L. Malec, L. Kaper, W. Ubachs, Phys. Rev. Lett. 106, 180802 (2011).

bA. Malec, R. Buning, M.T. Murphy, N. Milutinovic, S.L. Ellison, J.X. Prochaska, L. Kaper, J. Tumlinson, R.F. Carswell, W.

Ubachs, Mon. Not. Roy. Astron. Soc. 403, 1541 (2010).

cE.J. Salumbides, M.L. Niu, J. Bagdonaite, N. de Oliveira, D. Joyeux, L. Nahon, W. Ubachs, Phys. Rev. A 86, 022510 (2012).

dJ. Bagdonaite, P. Jansen, C. Henkel, H.L. Bethlem, K. Menten, W. Ubachs, Science 339, 46 (2013).